Visiting the Italian town of Pisa, young physics student Isaac Galilei drops an apple with a weight of 2.0 N from the city's famous leaning tower. The apple falls 56 m to the ground. We've just seen that the apple takes 3.4 s to reach the ground. so… while the apple falls downwards, how far does the Earth fall up towards the apple? (Ignore all effects except the Earth-apple interaction.) Should you need them, here are some data: The mass of the Earth is 5.97 x 10^24 . (You don't need to know that its radius is 6371 km and that G = 6.67384 × 10^ -11 m^3 kg^-1s^-2.
1
Expert's answer
2020-04-23T12:45:55-0400
"a_E=\frac{W}{M_E}"

"h_E=0.5a_Et^2=0.5t^2\frac{W}{M_E}"

"h_E=0.5(3.4)^2\frac{2}{5.97\cdot10^{24}}=1.9\cdot10^{-24}\ m"